### v3

# Streamlining Redundant Layers to Compress Large Language Models

Abstract
This paper introduces LLM-Streamline, a novel layer pruning approach for large language models. It is based on the observation that different layers have varying impacts on hidden states, enabling the identification of less important layers. LLM-Streamline comprises two parts: layer pruning, which removes consecutive layers with the lowest importance based on target sparsity, and layer replacement, where a lightweight network is trained to replace the pruned layers to mitigate performance loss. Additionally, a new metric called “stability” is proposed to address the limitations of accuracy in evaluating model compression. Experiments show that LLM-Streamline surpasses previous state-of-the-art pruning methods in both accuracy and stability.
Figure 1: The left figure illustrates the LLM-Streamline workflow, which includes layer pruning to remove consecutive layers and layer replacement where a lightweight network is trained to replace the pruned layers. The right figure presents the comparison results of LLM-Streamline with previous state-of-the-art (SOTA) pruning methods on 12 classification benchmarks (details in Section 4.2) by pruning 25% of the parameters on Llama2-7B. LLM-Streamline achieves 11.2% higher relative accuracy, which is the accuracy of the compressed model compared to the original model, than the previous SOTA.
1 Introduction
Large language models (LLMs) based on transformer architecturesvaswani2017attention have garnered significant attention and are employed across various domains and tasks. However, as LLMs increase in size, their hardware requirements escalate substantially, thereby constraining their applicability and impeding their deployment in real-world scenarios.
To reduce the hardware requirements for deploying LLMs, research efforts focus on developing compact models that maintain high performance, often through model compression techniqueszhu2023survey ; wang2024survey . These compact models significantly reduce memory consumption, computation time, and deployment costs. Current model compression techniques are broadly categorized into knowledge distillationhinton2015distilling ; gou2021knowledge ; li2022apidistill ; huang2022distill ; ho2022distill , model quantizationliu2021quan ; gholami2022survey ; dettmers2022int8 , and model pruninglouizos2017l0 ; chen2023lorashear ; frantar2023sparsegpt ; das2023beyond ; sun2023simple ; xia2023sheared . Quantization compresses the model by quantizing the weights to lower precision. Distillation achieves compression by transferring the capabilities of a larger teacher model to a smaller, more efficient student model. On the other hand, pruning compresses the model by eliminating unimportant parameters and modules.
Current pruning methods for LLMs primarily focus on pruning dense matricesashkboos2024slicegpt , attention headsmichel2019sixteen ; voita2019mha , filtersmccarley2019filter ; prasanna2020filter , and hidden dimensionsxia2023sheared ; van2023llm , with less emphasis on layer pruning. These methods often result in structural irregularities, making the pruned models difficult to store and deploy. In contrast, the layer pruning method, simply modifies the depth of the LLMs***The layers of LLMs are stored in data structures like nn.ModuleList, so layer pruning only requires removing elements from this list, making it easy to implement., is easier to implement. Therefore, exploring an effective layer-wise pruning method is crucial.
The effect of each layer in LLM can be viewed as a transformation of the input hidden state. We investigate the significance of different layers and the potential for pruning based on these transformations. As shown in Fig. 2, we observe that certain contiguous layers have minimal impact on the hidden state, indicating they are less important. This allows for efficient layer pruning by replacing these less important layers with lighter networks.
Based on the empirical observations described above, we propose LLM-Streamline, a straightforward and effective layer-wise pruning algorithm for LLMs. Our algorithm consists of two parts: layer pruning, in which we remove several consecutive layers with the lowest importance from the original model according to the target sparsity; and layer replacement, in which we train a lightweight network to replace the pruned layers to recover the performance degradation due to pruning. We experiment with various structures as lighter networks to replace the pruned layers, including feed-forward neural network (FFN), SwiGLU-based feed-forward neural network (SwiGLU), and transformer layer (Layer). Additionally, we find that removing layers without replacement significantly decreases model performance, highlighting the importance of layer replacement.
Moreover, we observe that the existing accuracy metric for evaluating model compression methods has limitations. For natural language understanding tasks involving multiple-choice classification, the compressed model might guess the correct answers for samples which the original model is uncertain about, leading to an overestimation of model compression performance. To address this, we propose a new metric called “stability” to emphasize the original model’s confidence and the consistency between the model before and after pruning.
Overall, this paper makes the following contributions:
  • We propose LLM-Streamline, a layer-wise pruning algorithm that removes less important layers by analyzing their transformations to hidden states in LLMs. To mitigate the potential performance loss caused by pruning, we innovatively employ a lightweight network to approximate the functionality of the pruned layers, thereby compensating for the loss incurred.
  • We propose a new metric called “stability”, which considers both the confidence of the original model in its answers and the consistency between the pruned and original model’s answers. The stability metric more accurately reflects the performance of the pruned model in classification tasks compared to the existing accuracy metric.
  • We conduct experiments on 12 well-known classification benchmarks and 3 generation benchmarks for LLMs. Our results show that for an LLM with 7B parameters and a 25% pruning rate, we can maintain 93% performance in classification tasks and 77% in generation tasks with only 30k data samples, outperforming existing SOTA pruning methods.
Note: As we were finalizing this work, some preprints men2024shortgpt ; gromov2024unreasonable (ShortGPT and UIDL) were posted, presenting a similar pruning method to ours. However, we have also proposed an additional layer replacement component.
Figure 2: The cosine similarity between the input and output hidden states of the layers in OPT-1.3B, OPT-2.7B, OPT-6.7B, and Llama2-7B.
2 LLM-Streamline
The workflow of the LLM-Streamline framework, shown in Fig. 1 (a), comprises two main steps: layer pruning and layer replacement. First, we prune redundant layers from the LLMs. Then, we train a lightweight network to replace the pruned layers and restore the model’s performance.
2.1 Layer Redundancy in LLMs
LLMs primarily utilize a transformer architecture, consisting of a series of transformer decoder layers. These layers adopt a residual structure, meaning the effect of each transformer layer can be viewed as a transformation of the input hidden state. Assuming that for the $\ell$-th layer $f$, the parameters are $\theta^{(\ell)}$ and the input hidden states are $\boldsymbol{x}^{(\ell)}$, the effect of layer $f$ can be formalized as
$\boldsymbol{x}^{(\ell+1)}=\boldsymbol{x}^{(\ell)}+f(\boldsymbol{x}^{(\ell)},\theta^{(\ell)}).$ (1)
In (1), the $\ell$-th layer $f$ contributes a transformation $f(\boldsymbol{x}^{(\ell)},\theta^{(\ell)})$ to the input $\boldsymbol{x}^{(\ell)}$. Therefore, we assess the importance of each layer in LLMs by evaluating its impact on the input hidden states. We use the cosine similarity $\cos{(\cdot,\cdot)}$ between inputs $\boldsymbol{x}^{(\ell)}$ and outputs $\boldsymbol{x}^{(\ell+1)}$ as a metric. Essentially, a higher cosine similarity between the inputs and outputs of a layer indicates lower importance, and vice versa. This interpretation arises from the observation that a high cosine similarity suggests the layer’s transformation is minimal, making it more amenable to pruning.
To measure the importance of different layers in LLMs, we randomly select samples from the pre-training data (details in Section 4.1). We then record the hidden states generated by the LLMs for these samples and compute the cosine similarity between the input and output hidden states of each layer. The computation of cosine similarity can be formalized as follows:
$\cos{(\boldsymbol{x}^{(\ell)},\boldsymbol{x}^{(\ell+1)})}=\mathbb{E}_{(\boldsymbol{x}^{(\ell)}_{i},\boldsymbol{x}^{(\ell+1)}_{i})\in\mathcal{D}}\left(\frac{1}{L}\sum_{j=1}^{L}\frac{\boldsymbol{x}^{(\ell)}_{i,j}\cdot\boldsymbol{x}^{(\ell+1)}_{i,j}}{\|\boldsymbol{x}^{(\ell)}_{i,j}\|\cdot\|\boldsymbol{x}^{(\ell+1)}_{i,j}\|}\right),$ (2)
where $\mathcal{D}$ denotes the recorded hidden states from different samples, $\boldsymbol{x}^{(\ell)}_{i}$, $\boldsymbol{x}^{(\ell+1)}_{i}\in R^{d\times L}$ denotes the inputs and outputs hidden states of the $i$-th sample respectively, $d$ denotes the hidden size and $L$ denotes the sequence length of each sample.
To mitigate the effects of model size and model structure, we conduct experiments on four models OPT-1.3B, OPT-2.7B, OPT-6.7Bzhang2022opt and Llama2-7Btouvron2023llama , across two structures and four sizes. The results, illustrated in Fig. 2, show that several consecutive layers in all models exhibit high cosine similarity, indicating a low level of importance.
2.2 Layer Pruning
As Fig. 2 shows, these least important layers are often contiguous. Then given $n$ determined by the target sparsity, we remove $n$ contiguous layers by finding the initial layer $\ell^{*}(n)$ with the highest cosine similarity for pruning, where $n$ is determined by the target sparsity:
$\ell^{*}(n)=\underset{\ell}{\arg\max}\,\cos(\boldsymbol{x}^{({\ell})},\boldsymbol{x}^{({\ell+n})}),$ (3)
where we randomly select samples from the pre-training data to compute the cosine similarity between $\boldsymbol{x}^{(\ell)}$ and $\boldsymbol{x}^{(\ell+n)}$, as outlined in Eq. (2).
2.3 Layer Replacement
After the layer pruning process, we aim to replace the pruned layer with a lightweight network that has fewer parameters. The rationale is that these layers contribute only minor transformations to the inputs. Therefore, we hypothesize that the cumulative effect of these layers can be approximated by a lightweight network. Specifically, after identifying the initial layer $\ell^{*}(n)$ for pruning, we use $(\boldsymbol{x}^{(\ell^{*})},\boldsymbol{x}^{(\ell^{*}+n)})$ as the training data to train the lightweight network using mean squared error (MSE) loss, which can be formalized as follows:
[…]
where $h$ denotes the lightweight network, $\mathcal{D}$ denotes the recorded hidden states of samples.
Discussion. The potential doubt is why to use layer replacement, instead of using common Parameter-Efficient Fine-Tuning (PEFT) methods such as LoRA hu2021lora and QLoRA dettmers2023qlora following layer pruning.
First, from the perspective of resource overhead, layer replacement is more adaptable to hardware resource constraints compared to other methods. Fine-tuning the model using the PEFT methods requires storing the model’s weights, activation values, and the optimizer state of the PEFT module in the GPU during training. In contrast, layer replacement involves two stages: dataset construction and model training. The first stage only requires storing the model’s weight and the forward computation overhead, while the second stage requires only the storing of a lightweight network. Therefore, layer replacement can also be implemented under hardware resource constraint conditions.
Second, layer replacement with a lightweight network, which substitutes for pruned layers, ensures that the pruned model more closely resembles the original model compared to models derived using PEFT methods. This similarity leads to a higher potential for performance recovery in the pruned model. This will be validated in our experiments.
Dataset #$\#{\rm TP}$$\#{\rm FN}$$\#{\rm FP}$$\#{\rm TN}$
C3 543 257 210 815
CHID 269 563 177 993
Race-M 380 95 129 832
Race-H 938 305 353 1902
Model ${\rm TP}$ ${\rm FN}$ ${\rm FP}$ ${\rm TN}$
Llama2-7B 1.12 0.87 0.94 1.02
w/ pruning 1.13 0.84 0.88 0.92
Table 1: (a) Number of samples in ${\rm TP},{\rm FN},{\rm FP}$, and ${\rm TN}$. (b) The PPL standard deviation results (×$\times 10^{-3}$) for Llama2-7B and its pruned version on Race-H.
3 Metrics for Evaluating Pruned Models
Accuracy is the most commonly used metric when evaluating LLMs in classification tasks. However, accuracy may overestimate the performance of the model after compression, since it does not take into account the consistency of the model’s answers before and after compression. In this section, we analyze such limitation and propose a novel metric for compressed models.
3.1 Shortcoming of Accuracy Metric
When evaluating the natural language understanding ability of LLMs, most existing benchmarks formalize the task as a classification task. The input to these tasks is a question accompanied by several choices, and the LLM is required to select the correct answer from these choices. Typically, each choice is concatenated with the question to form multiple sentences, and the perplexity (PPL) of each is computed. The choice corresponding to the sentence with the lowest PPL is selected as the answer. A classification task with $k$ choices and consisting of $N$ samples is denoted as $\mathcal{T}=\{\left(x_{i},c_{i,1},c_{i,2},...,c_{i,k},y_{i}\right)\}_{i=1}^{N}$, where $x_{i}$ represents the question in the $i$-th sample, $c_{i,j}$ represents the $j$-th choices, and $y_{i}$ represents the correct choice. Additionally, we define $\mathcal{M}$ to denote the original LLM, $\bar{\mathcal{M}}$ to denote the compressed LLM, and $\hat{y}(\mathcal{M})$ to denote the choice predicted by the model $\mathcal{M}$. We initially evaluate the performance of the model pruned by the method described in Sec 2.2 and observe the performance of the pruned model has been improved on some classification tasks. To further investigate this phenomenon, we analyze the experimental results using the confusion matrix li2024evaluating . Specifically, we count the number of samples and average standard deviation (std) for the PPL of the samples for each term of the confusion matrix. The calculation of the std for the PPL of the $i$-th sample is defined as follows:
$\small\mathrm{PPL}_{i,j}=\mathrm{PPL}(\mathcal{M}(x_{i},c_{i,j})),{\mathrm{PPL}}_{i}=\frac{\sum_{j=1}^{k}\mathrm{PPL}_{i,j}}{k},\mathrm{std}_{i}=\sqrt{\frac{\sum_{j=1}^{k}(\mathrm{PPL}_{i,j}-{\mathrm{PPL}}_{i})^{2}}{k-1}},$ (5)
where ${\rm PPL}_{i,j}$ denotes the PPL for the sentence created by question $x_{i}$ and choice $c_{i,j}$ of the model before pruning, ${\rm std}_{i}$ denotes the std for PPL of the $i$-th sample. A higher $\mathrm{std}_{i}$ value indicates the LLM exhibits greater confidence in answering the question $x_{i}$.
Each term of the confusion matrix is defined as follows,
  • ${\rm TP}\left[\hat{y}_{i}({\mathcal{M}})=y_{i}\wedge\hat{y}_{i}(\bar{\mathcal{M}})=y_{i}\right]$ is a set of samples where the model answers correctly both before and after pruning.
  • ${\rm FN}\left[\hat{y}_{i}({\mathcal{M}})=y_{i}\wedge\hat{y}_{i}(\bar{\mathcal{M}})\neq y_{i}\right]$ is a set of samples where the model answers correctly before pruning but incorrectly after pruning.
  • ${\rm FP}\left[\hat{y}_{i}({\mathcal{M}})\neq y_{i}\wedge\hat{y}_{i}(\bar{\mathcal{M}})=y_{i}\right]$ is a set of samples where the model answers incorrectly before pruning but correctly after pruning.
  • ${\rm TN}\left[\hat{y}_{i}({\mathcal{M}})\neq y_{i}\wedge\hat{y}_{i}(\bar{\mathcal{M}})\neq y_{i}\right]$ is a set of samples where the model answers incorrectly both before and after pruning.
Table 1 presents the counts of samples in TP, FN, FP, TN, and also the PPL standard deviation in several datasets. We can observe that samples in ${\rm FP}$ constitute a considerable proportion of the total samples, and the std for ${\rm TP}$ and ${\rm TN}$ is significantly higher than that for ${\rm FN}$ and ${\rm FP}$. This indicates that the model is more uncertain about the ${\rm FN}$ and ${\rm FP}$ samples, implying that the model may guess the correct answer for a significant portion of these samples after pruning. This phenomenon suggests that the accuracy metric may overestimate the performance of the compressed model.
3.2 Stability Metric
We propose a novel metric “stability” to evaluate the performance of LLMs after pruning, i.e.,
$\mathrm{Stability}(\mathcal{M},\bar{\mathcal{M}})=\frac{\sum_{i=1}^{N}\left(\exp{(\mathrm{std}_{i})}\cdot\mathbbm{1}_{\left[i\in{\rm TP}\cup{\rm TN}\right]}\right)}{\sum_{i=1}^{N}\exp{(\mathrm{std}_{j}})},$ (6)
where the identifier $\mathbbm{1}_{\left[i\in{\rm TP}\cup{\rm TN}\right]}$ is used to indicate whether the $i$-th sample belongs to ${\rm TP}$ and ${\rm TN}$. We use ${\rm std}_{i}$ as the weight of the $i$-th sample. Because the ${\rm std}$ of different samples varies significantly. to mitigate the influence of samples with excessively large standard deviations, we apply the $\exp$ function to moderate the weight differences among samples. Different from accuracy, stability focuses on the model’s confidence in its answers and the consistency between the model before and after pruning on a task, aligning more closely with the goal of model pruning, i.e., ensuring the pruned model remains as similar as possible to the original model.
4 Experiments
In this section, we first compare our proposed method, LLM-Streamline, with several popular structured pruning methods to demonstrate the effectiveness of our method (4.4). We also highlight the necessity of the lightweight network by ignoring it. After that, we analyze the impact of different sizes and structures of lightweight networks on the model’s performance (4.5) and evaluate the performance under various pruning ratios (4.6). Finally, we compare the proposed layer replacement with the well-known PEFT method, LoRA hu2021lora (4.7), demonstrating that layer replacement provides superior performance and reduced memory overhead.
4.1 Setup
To validate the effectiveness of our proposed model, we conduct experiments on popular open-source LLMs, including OPT-6.7Bzhang2022opt and Llama2-7B touvron2023llama . Following previous work, we perform experiments with a 25% compression rate and extract data from SlimPajama for layer pruning and layer replacement, where SlimPajama is a dataset containing data from different domains for pre-training LLM. Sheared LLaMa xia2023sheared finds that the performance degradation of pruned models varies across different domains, and proposes determining the distribution of data from different domains based on the degree of performance degradation. Therefore, we randomly sample the data based on the distribution used by Sheared LLaMa xia2023sheared , finally constructing the dataset containing 30,000 pieces of data. We randomly select 500 samples from this dataset and input them into LLMs, generating Fig. 2, and use these 500 data samples for layer pruning. All 30,000 pieces of data are used to train the lightweight network. We use two types of lightweight networks: Feed-Forward Neural Network (FFN) and Transformer Layer. The FFN is randomly initialized, while the Transformer Layer inherits the parameters of the first pruned layer. When training the FFN and Transformer layers, different learning rates are used: $10^{-3}$ for the FFN and $10^{-5}$ for the Transformer layer. Both are trained for $20$ epochs with a batch size of $32$. On an A800 GPU, the training process takes approximately $5$ hours to complete. Further experimental details are available in the Appendix B.1.
4.2 Benchmark
To comprehensively assess the changes in LLM capabilities before and after pruning, we follow the evaluation method of LaCo yang2024laco . We use 12 natural language understanding and question-answering benchmarks for evaluation: CMNLI xu2020clue , HellaSwag(HeSw) zellers2019hellaswag , PIQA bisk2020piqa ,CHID zheng2019chid , WSC levesque2012winograd ,CommonSenseQA(CoQA) talmor2018commonsenseqa , BoolQ clark2019boolq ,MMLU hendrycks2020mmlu , CMMLU li2023cmmlu ,Race-High/Middle lai2017race , C3 sun2020C3 . The tasks in these benchmarks are formalized as classification tasks, so we refer to these benchmarks as classification benchmarks. For these benchmarks, we use both accuracy and stability as metrics for evaluating the models. Additionally, we include 3 benchmarks: XSum narayan2018xsum , GSM8K cobbe2021gsm8k and StrategyQA geva2021did , to demonstrate the LLM’s performance on generation tasks after pruning. We refer to these tasks as generation benchmarks. Following the evaluation framework of OpenCompasscontributors2023opencompass , we use accuracy to evaluate StrategyQA and GSM8K, and use ROUGE1 to evaluate Xsum.
4.3 Baseline
We compare several structured pruning methods that prune the attention head, the filters of the FFN layer, and the hidden dimension. We also compare existing layer-pruning methods.
LLM-Prunerma2023llmpruner prunes attention heads, FFN layer filters, and hidden dimensions by using gradients and activations to estimate the importance of these modules.
SliceGPTashkboos2024slicegpt prunes hidden dimensions. It inserts dimensionality reduction matrices into the model and employs Principal Component Analysis (PCA) to initialize and eliminate unimportant components within these matrices. The reduced-dimensional matrices are subsequently merged with the model’s original weight matrices, thereby reducing the model’s size.
LaCoyang2024laco prunes layers by dividing the layers into groups, each consisting of multiple consecutive layers, and compress them separately, whereas our method simply compresses a piece of consecutive layers. LaCo merges consecutive layers by averaging their parameters whereas we train an additional lightweight network to replace these layers.
LLM Method Ratio Benchmarks Ave. Per.
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Llama2-7B Dense 0.00% 43.8 33.0 41.6 70.8 37.5 66.7 71.3 78.1 33.1 35.5 46.8 31.8 49.2 100.0
LLMPruner 24.8% 29.7 33.4 28.4 58.7 40.4 48.5 54.6 72.0 22.9 22.0 25.3 25.0 38.4 78.0
SliceGPT 25.4% 31.5 31.6 18.5 59.9 43.3 49.6 47.5 68.3 27.0 29.4 28.8 24.8 38.4 78.0
LaCo* 27.0% 39.7 34.4 36.1 64.1 40.4 45.7 55.7 69.8 23.6 22.6 26.5 25.2 40.3 81.9
Ours (None) 24.0% 40.2 34.4 21.5 67.3 40.4 51.7 59.7 69.0 35.2 34.7 44.6 28.9 44.0 89.4
Ours (FFN) 25.0% 40.7 33.0 22.8 65.9 38.5 60.6 61.2 71.2 38.0 38.7 47.0 31.7 45.8 93.1
Ours (Layer) 24.0% 43.3 33.0 24.1 67.5 36.5 59.2 61.1 71.5 34.8 37.0 45.5 29.4 45.2 91.9
OPT-6.7B Dense 0.00% 38.7 32.9 21.6 64.6 41.4 54.8 63.3 75.4 25.1 25.4 24.7 25.5 41.1 100
SliceGPT 25.6% 40.0 31.2 19.5 37.9 36.5 38.2 45.6 65.8 25.8 26.0 25.8 24.8 34.8 84.7
Ours (None) 24.0% 27.6 32.5 12.7 44.8 36.5 20.6 26.5 52.1 22.1 22.4 23.6 25.2 28.9 70.3
Ours (FFN) 25.0% 37.6 32.1 18.7 63.7 37.5 41.8 55.9 73.2 22.7 22.2 24.4 24.9 37.9 92.2
Ours (Layer) 24.0% 36.4 32.0 18.9 62.4 38.5 45.1 54.3 74.0 23.6 24.2 24.3 25.2 38.2 92.9
Table 2: Accuracy of pruning methods on classification benchmarks. “*” indicates that we refer to the results in the original paper. ShortGPT and UIDL men2024shortgpt ; gromov2024unreasonable can be considered as a special case of our approach, i.e., Ours (None). Preformance (Pre.) represents the percentage of the original model’s performance retained by the pruning method.
LLM Method Ratio Benchmarks Ave.
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Llama2-7B LLMPruner 24.8% 72.8 94.0 74.1 70.8 87.5 71.0 79.9 86.8 52.4 55.2 53.3 65.9 72.0
SliceGPT 25.4% 53.2 35.4 53.3 77.1 80.8 75.3 71.6 78.7 90.7 85.3 60.3 56.7 68.2
Ours (None) 24.0% 76.6 38.7 65.3 81.4 87.5 74.7 80.7 81.0 73.7 67.9 80.1 70.8 73.2
Ours (FFN) 25.0% 79.8 100 64.1 83.1 93.3 80.7 84.7 84.6 85.1 79.0 87.5 82.5 83.7
Ours (Layer) 24.0% 79.8 100 64.4 86.3 95.2 81.7 85.3 85.6 81.8 79.0 82.4 71.0 82.7
OPT-6.7B SliceGPT 25.6% 66.4 39.8 73.1 30.1 79.4 75.2 73.7 82.5 72.5 69.1 68.9 67.2 66.5
Ours (None) 24.0% 56.2 62.5 71.4 41.5 87.5 48.6 50.5 57.1 63.4 62.2 62.3 62.7 60.5
Ours (FFN) 25.0% 74.1 36.1 77.4 74.2 90.4 82.4 88.4 92.0 68.3 63.9 74.7 65.0 73.9
Ours (Layer) 24.0% 72.1 35.5 76.1 72.1 91.4 83.6 87.0 91.6 72.9 71.2 78.1 70.1 75.1
Table 3: Stability of pruning methods on classification benchmarks. The stability of the original model is 1.0, because stability is measured by comparing the prediction results of the original model.
LLM Method Ratio Benchmarks Ave. Per.
Xsum GSM8K StrategyQA
Llama2-7B Dense 0.00% 19.4 16.5 60.2 32.0 100.0
LLMPruner 24.8% 16.4 0.61 44.2 20.4 63.8
SliceGPT 25.4% 12.4 3.34 45.7 20.5 64.1
LaCo* 27.1% 15.6 - - - -
Ours (None) 24.0% 14.8 1.97 41.8 19.5 60.9
Ours (FFN) 25.0% 18.6 2.16 46.5 22.4 70.0
Ours (Layer) 24.0% 20.2 1.82 52.1 24.7 77.2
OPT-6.7B Dense 0.00% 13.4 2.2 54.3 23.3 100.0
SliceGPT 25.6% 14.9 2.5 40.8 19.4 83.3
Ours (None) 24.0% 4.9 0 0 1.6 6.87
Ours (FFN) 25.0% 14.8 0.8 43.6 19.7 84.5
Ours (Layer) 24.0% 18.4 2.5 44.4 21.8 93.6
Table 4: Evaluations on generation benchmarks. “*” indicates that we refer to the results in the original paper.
4.4 Main Results
We present accuracy and stability for different methods on the classification benchmarks in Table 2 and Table 3, respectively, and Table 4 for the generative benchmarks. The results demonstrate that our proposed LLM-Streamline consistently outperforms the baseline methods, effectively retaining the majority of the capabilities of LLMs across various tasks. Specifically, in classification tasks, on Llama2-7B, LLM-Streamline surpasses SliceGPT by 7% in terms of accuracy and 15% in terms of stability. LLM-Streamline also surpasses LaCo by 5% in terms of accuracy. For generation tasks, LLM-Streamline retains nearly 77% of Llama2-7B’s capabilities and 94% of OPT-6.7B’s capabilities, significantly outperforming other pruning methods. We find that almost all of the pruning methods fail on the GSM8K dataset, which reflects the difficulty of math word problems.
Additionally, comparing the average stability (Ave.) in Table 3 with the retrained performance (Pre.) in Table 2 across different pruning methods reveals that stability is often much lower than accuracy. We also observe that accuracy on Race-M and Race-H even increases after model pruning. This suggests that pruned models tend to make correct guesses on some classification questions that they are uncertain, highlighting the limitations of accuracy as a sole measure of pruning method performance. We also conduct experiments on smaller models, such as OPT-1.3B and OPT-2.7B. Details can be found in Appendix  C.1.
Layer-Random Layer-First Layer-Last Layer-Avg FFN FFN SwiGLU SwiGLU
Accuracy 45.1 45.2 45.6 44.4 46.0 45.8 43.8 44.2
Stability 81.2 82.7 81.9 79.2 80.7 83.7 82.6 83.3
[…]
Figure 3: Validation loss curves during training of (a) FFN and SwiGLU; (b) Transformer layer.
Figure 4: (a) Stability of the pruned Llama-7B at pruning ratios of about 2%, 11%, 20%, and 25%. Since stability is calculated solely for compressed model compared to the original model, we could not compute the stability for TinyLlama-1.1B and OpenLlama-3B. (b) Accuracy of the pruned LLama-7B at the same pruning ratios. Metrics are averaged across the classification benchmarks.
4.5 Impact of Different Lightweight Networks
To investigate the impact of the model size, architecture, and initialization methods of lightweight networks on the pruned model’s performance, we conduct experiments on Llama2-7B using the different architecture as lightweight network, including Feed-Forward Neural Network (FFN), SwiGLU-based Feed-Forward Neural Network (SwiGLU), and Transformer Layer (Layer). We also explore various initialization methods for the Transformer Layer, including random initialization (Layer-Random), inheritance of the first pruned layer (Layer-First), inheritance of the last pruned layer (Layer-Last), and averaging the pruned layers (Layer-Avg). The average accuracy and stability metrics across all the classification benchmarks are presented in Table 5, with detailed results on each benchmark available in Appendix C.2. The results show that FFN achieves the best results. Meanwhile, for the Transformer Layer, inheritance of the pruned first layer yields the best results. In addition, We plot the validation loss curves during the training process for different lightweight networks, as shown in Fig. 3. We can observe that FFN and SwiGLU have already converged by the 10th epoch, whereas the loss of Transformer Layer is still decreasing. This indicates that the Transformer layer still has potential, and further training could yield better results, but this would require more computing resources.
4.6 Impact of Different Pruning Ratios
To verify the performance of the model at other pruning ratios, in addition to the about 25% pruning ratio, we also evaluate our method’s performance at other pruning ratios on Llama2-7B, with pruning ratios 2%, 11%, and 20%. The average stability and accuracy metrics across all the classification benchmarks are shown in Fig. 4, with details on each benchmark presented in Appendix C.3. By comparing the performance of Llama2-7B, TinyLlama-1.1B, OpenLlama-3B-v2, and the model at different pruning ratios, we approximate that the performance of the pruned model is linearly correlated with the number of parameters, demonstrating the effectiveness of our method.
4.7 Comparison of Layer Replacement and LoRA
We compare the performance of layer replacement with LoRA. Since layer replacement is trained based on hidden states with a different training objective than LoRA, we additionally train one epoch using the language model loss for layer replacement when comparing it with LoRA. The training details can be found in the Appendix B.2. For layer replacement, we freeze the original model’s weights and only train the lightweight network. For LoRA, we set the rank to 384 to ensure that the number of parameters involved in training is approximately the same as that of lightweight networks. We randomly extract 30,000 samples from SlimPajama-6B as training data. The average accuracy and stability across all the classification benchmarks are presented in Table 6, with detailed results for each benchmark available in Appendix C.4. The results show that layer replacement outperforms LoRA in both accuracy and stability. In addition to the impact on performance, the post-training process for layer replacement also demands significantly less GPU memory than LoRA requires.
Layer-First Layer-Last Layer-Avg FFN FFN SwiGLU SwiGLU LoRA
Accuracy 46.7 46.8 46.2 45.8 46.3 44.4 45.5 44.6
Stability 85.7 85.6 83.9 83.4 85.2 84.7 84.7 82.6
GPU Memory (G) 27.8 27.8 27.8 25.6 27.0 25.3 26.4 47.6
[…]
5 Related Work
Model pruning can be categorized into unstructured pruning and structured pruning. Unstructured pruning compresses the model by converting its dense matrices into sparse matrices, retaining the dense matrix, and is more hardware-friendly louizos2017l0 ; frantar2023sparsegpt ; das2023beyond ; sun2023simple . Structured pruning maintains performance at high compression rates but requires specialized sparse operators, limiting its applicability chen2023lorashear ; xia2023sheared ; ashkboos2024slicegpt ; ma2023llmpruner .
LLM-Prunerma2023llmpruner is a structured pruning method that uses gradients and activations to estimate the importance of groups of modules in LLMs and prunes the less important ones. SliceGPTashkboos2024slicegpt , another structured pruning method, inserts dimensionality reduction matrices into LLMs and uses PCA to initialize and eliminate unimportant components within these matrices. LaCoyang2024laco prunes layers by merging similar layers. Unlike LaCo, our proposed method, LLM-Streamline, uses a lightweight network to compensate for the loss caused by the pruned layers instead of merging them.
6 Conclusion
This work introduces a layer pruning-and-replacement algorithm for LLMs. We identify shortcomings in the existing accuracy metric and propose a new metric called “stability” for model compression. Extensive experiments show that this layer replacement method using a lightweight network outperforms previous state-of-the-art pruning methods.
Appendix A Data Distribution
We extract the training data from different domains based on the data distribution strategy proposed in Sheared-LLaMa [17]. The detailed data distribution is shown in Table 7.
CC GitHub Book StackExchange Wiki ArXiv C4
SlimPajama-6B 54.1% 4.2% 3.7% 2.8% 3.1% 3.4% 28.7%
Ours 36.1% 0.8% 9.1% 1.0% 3.1% 0.7% 49.2%
Table 7: The proportion of different domains randomly selected from the SlimPajama-6B dataset.
Appendix B Training Implementation Details
B.1 Lightweight Network Training Details
For both the FFN structure and the SwiGLU structure, the learning rate is set to 1e-3 and the weight decay is 1e-4. For the Transformer layer, the learning rate is set to 1e-5 and the weight decay is 1e-3. The model is trained using a batch size of 32 over 20 epochs. On a single A800 GPU, the training duration for the lightweight network is approximate 5 hours (for the Transformer layer).
B.2 Post Training Details
For layer replacement, in order to have a fairer comparison with LoRA, we conduct one epoch of post-training with a learning rate of 5e-5, a weight decay of 1e-3, and a batch size of 32. This process takes less than an hour on a single A800 GPU. For LoRA, the model is trained over 10 epochs with a learning rate of 5e-5, a weight decay of 1e-3, and a batch size of 32, requiring approximate 10 hours on a single A800 GPU.
Appendix C Detailed Experimental Results
C.1 Experimental Results of OPT-1.3B and OPT-2.7B
We also conduct experiments on small models (OPT-1.3B and OPT-2.7B). The experimental results are shown in Table 8. The results indicate that our proposed LLM-Streamline is superior to the previous SOTA method, SliceGPT, across different pruning rates.
LLM Method Ratio Benchmarks Ave. Per.
PIQA WinoGrande HellaSwag ARC-easy ARC-challenge OpenBookQA
OPT-1.3B Dense 0.00% 72.4 59.3 53.7 51.0 29.5 23.4 48.2 100.0
SliceGPT 18.1% 67.6 53.6 35.7 51.1 23.1 20.2 41.9 86.9
Ours(None) 19.4% 57.2 51.7 29.1 32.5 22.7 13.2 34.4 71.4
Ours(FFN) 18.1% 68.8 58.4 39.1 54.3 23.3 23.3 44.5 92.3
Dense 0.00% 72.4 59.3 53.7 51.0 29.5 23.4 48.2 100.0
SliceGPT 25.8% 65.5 52.8 34.2 48.8 24.4 17.0 40.5 84.0
Ours(None) 27.1% 52.2 51.1 25.7 26.6 20.5 14.0 31.7 65.8
Ours(FFN) 25.8% 66.4 56.0 36.8 51.6 22.2 21.0 42.3 87.8
Dense 0.00% 72.4 59.3 53.7 51.0 29.5 23.4 48.2 100.0
SliceGPT 33.6% 62.4 52.6 32.2 45.4 23.1 16.6 38.7 80.3
Ours(None) 34.8% 50.5 51.5 25.8 26.2 20.3 14.6 31.5 65.4
Ours(FFN) 33.6% 62.9 52.1 33.9 48.3 20.8 20.6 39.8 82.6
OPT-2.7B Dense 0.00% 73.8 61.0 45.9 60.9 26.8 25.0 48.9 100.0
SliceGPT 16.8% 69.6 56.3 40.4 56.2 27.5 20.2 45.0 92.0
Ours(None) 17.8% 61.2 54.1 33.8 41.2 24.1 15.8 38.4 78.5
Ours(FFN) 16.8% 70.7 60.4 42.9 57.8 25.3 24.4 46.9 95.9
Dense 0.00% 73.8 61.0 45.9 60.9 26.8 25.0 48.9 100.0
SliceGPT 25.7% 69.1 55.0 37.9 53.9 26.7 18.2 43.5 89.0
Ours(None) 26.7% 59.7 53.4 33.5 38.1 24.3 15.4 37.4 76.5
Ours(FFN) 25.7% 67.0 59.5 40.3 54.6 24.7 22.2 44.7 91.4
Dense 0.00% 73.8 61.0 45.9 60.9 26.8 25.0 48.9 100.0
SliceGPT 34.6% 64.8 54.1 35.6 50.0 26.5 18.0 41.5 84.9
Ours(None) 35.6% 56.6 52.9 31.5 37.6 24.1 14.9 36.3 74.2
Ours(FFN) 34.6% 65.3 55.3 36.3 51.4 24.5 21.0 42.3 86.5
Table 8: Accuracy of different pruning methods by pruning OPT-1.3B and OPT-2.7B.
C.2 Detailed Results of Different Lightweight Networks
The detailed results of accuracy and stability of different lightweight networks on different classification benchmarks are shown in Table 9 and Table 10. We can observe that FFN achieves the best results. Meanwhile, for the Transformer Layer, inheritance of the pruned first layer yields the best results.
LLM Method Ratio Benchmarks Ave. Per.
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Llama2-7B Dense 0.00% 43.8 33.0 41.6 70.8 37.5 66.7 71.3 78.1 33.1 35.5 46.8 31.8 49.2 100.0
Layer-Random 24.0% 42.5 33.0 27.0 65.9 36.5 58.0 58.8 70.2 35.0 36.0 46.3 31.4 45.1 91.7
Layer-First 24.0% 43.3 33.0 24.1 67.5 36.5 59.2 61.1 71.5 34.8 37.0 45.5 29.4 45.2 91.9
Layer-Last 24.0% 43.5 33.0 29.0 64.5 41.4 56.8 61.5 71.6 34.5 35.0 46.0 30.8 45.6 92.7
Layer-Avg 24.0% 42.1 33.0 26.7 66.3 36.5 57.7 59.4 69.9 34.3 34.5 43.3 28.5 44.4 90.2
FFN 26.0% 40.6 33.0 24.2 67.5 36.5 58.4 59.5 71.4 41.9 41.4 46.3 30.8 46.0 93.5
FFN 25.0% 40.7 33.0 22.8 65.9 38.5 60.6 61.2 71.2 38.0 38.7 47.0 31.7 45.8 93.1
SwiGLU 26.0% 41.9 33.0 24.3 68.5 36.5 55.8 57.9 69.6 29.9 33.3 43.4 31.6 43.8 89.0
SwiGLU 25.0% 40.9 33.0 22.1 67.0 36.5 56.9 59.1 70.0 33.8 35.0 45.6 30.8 44.2 89.8
[…]
LLM Method Ratio Benchmarks Ave.
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Llama2-7B Layer-Random 24.0% 79.9 100 68.2 83.8 95.2 77.4 82.0 83.2 69.9 71.9 85.1 77.4 81.2
Layer-First 24.0% 79.8 100 64.4 86.3 95.2 81.7 85.3 85.6 81.8 79.0 82.4 71.0 82.7
Layer-Last 24.0% 81.0 100 73.5 84.9 82.7 81.3 85.8 85.4 82.3 76.3 83.0 66.9 81.9
Layer-Avg 24.0% 80.7 100 68.5 84.4 95.2 78.9 82.8 82.5 73.0 70.7 74.6 58.8 79.2
[…]
FFN 25.0% 79.8 100 64.1 83.1 93.3 80.7 84.7 84.6 85.1 79.0 87.5 82.5 83.7
SwiGLU 26.0% 78.5 100 64.5 78.9 95.2 75.4 80.9 82.1 89.3 87.1 80.7 78.4 82.6
SwiGLU 25.0% 78.9 100 63.5 84.4 95.2 77.0 82.2 82.3 85.7 82.4 87.9 79.7 83.3
[…]
C.3 Detailed Results of Different Pruning Ratio
The detailed results of accuracy and stability on LLMs under different pruning ratios on differnet classification benchmarks are shown in Table 11 and Table 12. The experiment results show that the performance of the pruned model is linearly correlated with the number of parameters, demonstrating the effectiveness of our method.
LLM Method Ratio Benchmarks Ave. Per.
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Llama2-7B Dense 0.00% 43.8 33.0 41.6 70.8 37.5 66.7 71.3 78.1 33.1 35.5 46.8 31.8 49.2 100.0
[…]
TinyLlama-1.1B Dense 0.00% 38.3 34.6 30.4 56.4 47.1 48.8 54.5 71.3 24.1 25.8 25.8 25.0 40.2 100.0
OpenLlama-3B-v2 Dense 0.00% 43.0 33.0 31.1 60.6 37.5 58.7 65.3 77.0 25.1 26.9 27.0 25.3 42.5 100.0
[…]
LLM Method Ratio Benchmarks Ave.
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
[…]
C.4 Detailed Results of Layer Replacement and LoRA
The detailed results of accuracy and stability of different layer replacement strategies and LoRA on different benchmarks are shown in Table 13 and Table 14. The results show that layer replacement outperforms LoRA in both accuracy and stability.
LLM Method Ratio Benchmarks Ave. Per.
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Llama2-7B Dense 0.00% 43.8 33.0 41.6 70.8 37.5 66.7 71.3 78.1 33.1 35.5 46.8 31.8 49.2 100.0
Layer-First 24.0% 43.9 33.0 29.8 70.8 36.5 59.6 64.3 73.4 36.6 37.4 44.9 30.0 46.7 94.9
Layer-Last 24.0% 44.9 33.0 29.4 69.2 36.5 58.9 63.5 74.1 37.8 37.8 46.5 30.4 46.8 95.1
Layer-Avg 24.0% 43.9 33.0 30.1 67.5 36.5 58.3 62.5 72.3 36.6 36.1 46.2 31.9 46.2 93.9
[…]
FFN 25.0% 43.8 33.0 27.0 68.7 36.5 60.7 63.5 72.4 37.4 35.4 45.3 31.5 46.3 94.1
SwiGLU 26.0% 44.0 33.0 27.9 61.2 36.5 57.2 61.7 71.2 30.3 32.9 45.0 31.9 44.4 90.2
SwiGLU 25.0% 43.2 33.0 27.1 67.0 36.5 58.2 62.1 71.2 35.1 35.7 45.8 30.8 45.5 92.5
LoRA 24.0% 43.7 33.0 28.6 62.5 36.5 57.4 62.5 71.7 30.9 32.9 43.7 31.2 44.6 90.7
[…]
LLM Method Ratio Benchmarks Ave.
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Llama2-7B Layer-First 24.0% 82.7 100 74.3 84.7 95.2 85.7 89.1 88.6 84.1 82.2 84.5 77.2 85.7
Layer-Last 24.0% 83.2 100 73.8 87.9 95.2 89.5 85.8 88.7 84.7 82.1 83.1 72.9 85.6
Layer-Avg 24.0% 81.0 100 72.3 67.0 95.2 84.1 87.4 86.6 87.3 82.2 90.5 73.7 83.9
[…]
FFN 25.0% 80.6 100 72.0 83.5 95.2 85.4 87.7 87.2 84.5 81.0 85.4 79.3 85.2
SwiGLU 26.0% 80.0 100 71.4 63.1 95.2 80.5 86.1 85.1 90.4 87.7 89.2 87.7 84.7
SwiGLU 25.0% 81.6 100 72.9 67.3 95.2 81.8 86.5 85.3 87.7 84.8 90.0 83.3 84.7
LoRA 24.0% 82.7 100 73.6 57.1 95.2 83.6 86.8 85.4 86.4 83.3 81.9 75.1 82.6
[…]